FedEx Corp. stock ended the previous year at $104.59 per share. It paid a $0.20 per share dividend last year. It ended last year at $107.89. If you owned 310 shares of FedEx, what was your dollar return and percent return? (Round your percent return answer to 2 decimal places.) Dollar return Percent return 0 0
Answer #1

Dollar return=(End price-Beginning price+Dividends)*310 shares

=(107.89-104.59+0.2)*310 shares=$1085

Percent return=Dollar return/Beginning price

=1085/(104.59*310)

=1085/32422.9=3.35%(Approx).